OpenStudy (lgbasallote):

i need to put this in the form of linear differential equation (the one with the D's) \[y = 2e^x \cos 3x\]

OpenStudy (lgbasallote):

i think \[m = 1 \pm 3i\]

OpenStudy (anonymous):

where is ur equation? lol ; i cant see

OpenStudy (lgbasallote):

\[y = 2e^x \cos x\] that's the equation lol

OpenStudy (lgbasallote):

i need to make it look like (D^2 -2D +10)y= 0

OpenStudy (anonymous):

\( m=1\pm 3i \) and \( (m-1)^2=-9 \) so \( m^2-2m+10=0 \)

OpenStudy (lgbasallote):

what did you do?

OpenStudy (lgbasallote):

ohh you transposed

OpenStudy (lgbasallote):

then squared huh

OpenStudy (anonymous):

yep

OpenStudy (ujjwal):

I have a question: what is 'm'?

OpenStudy (ujjwal):

@mukushla

OpenStudy (ujjwal):

A link would be helpful too if you are too bored to type a reply!